Session: Nov 2002 (session 3)
Wavelength: 18/21 cm


To give feedback on this experiment, click here

Experiment Feedback

Jodrell_Bank (Alastair Gunn at Tue Nov 19 07:19:26 2002) Success :
Did not use new ONOFF program due to conflicting logic with our local control software. Instead took calibration data manually.

Cambridge (Alastair Gunn at Tue Nov 19 07:19:57 2002) Success :
No TSYS information available through the MERLIN microwave link.

Torun (K.Borkowski at Tue Nov 19 10:26:37 2002) Success :
We did not update the .rxg file due to problems with the GNPLT software. No other known problems.

Urumqi (zhoujianjun at Wed Nov 20 02:33:20 2002) Minor failures : FS
the updated fs system has some problems in *.rxg file. it can not measures LCP and RCP simultaneously, so we measure the LCP only. we'll send you the log file of RCP.

Medicina (Orlati at Mon Nov 25 13:53:41 2002) Success :

Shanghai (Huang Xinyong at Tue Nov 26 01:13:47 2002) Success :
We done it on Nov.25, because of the RCP receiver was not stability and the weather was bad on Nov.19.

Effelsberg (A. Kraus at Tue Nov 26 10:10:12 2002) Success :

Onsala (John Conway at Tue Dec 3 18:01:10 2002) Failure :
No observations performed due to a telescope control problem.Onoff measuremenst were however made before and after the allocated slot.

Hartebeesthoek (Jonathan Quick at Wed Dec 11 12:03:18 2002) Failure :
Did not observe on correct date due to conflicting geodetic experiment but did collect data before and after.